OpenStudy (anonymous):

Find the lateral area for the cylinder with the given measurement. r = 6, h = 20

OpenStudy (theviper):

Lateral surface area of a Cylinder:- 

OpenStudy (theviper):

Therefore, area is\[2\times \frac{22}{7}\times6\times20\]

OpenStudy (theviper):

\[=\frac{5280}{7}=754.29\]

OpenStudy (anonymous):

My options were: a) 120(pi) b) 240(pi) c) 312(pi)

OpenStudy (theviper):

Ok!wait a minute

OpenStudy (theviper):

2*pi*6*20=240pi Can't u do that??

OpenStudy (anonymous):

That works. Thank You:)

OpenStudy (theviper):

yw:)